# A stochastic Gordon-Loeb model for optimal cybersecurity investment under clustered attacks

### v1

@@ Abstract
Keywords:
Cyber-risk; security breach; benefit-cost analysis; Hawkes process; stochastic optimal control.
@@ 1 Introduction
The problem of optimal cybersecurity investment has been first addressed in the seminal work of Gordon and Loeb [3]. In their model, reviewed in Section 2.1 below, the decision maker can reduce the vulnerability to cyberattacks by investing in cybersecurity. The optimal expenditure in cybersecurity is determined by maximizing the expected net benefit of reducing the breach probability. The Gordon-Loeb model laid the foundations for a rigorous quantitative analysis of cybersecurity investments and has been the subject of numerous extensions and generalizations: we mention here only some studies that are closely related to our context, referring to [4] for a comprehensive overview. The key ingredient of the Gordon-Loeb model is represented by the security breach probability function (see Section 2.1), which has been further analyzed in [5] and [6]. The risk-neutral assumption of [3] has been relaxed to accommodate risk-averse preferences in [7].
@@
An effective cybersecurity investment policy must be adaptive and evolve in response to changing threat environments. As noted by [11], a key feature of cyber-risk is its dynamic nature, due to the rapid technological transformation and the evolution of threat actors. Similarly, [12] emphasize the need for adaptable and responsive cybersecurity policies in order to face the challenge of dynamic cyberattacks. The framework of [10] is based on a deterministic model and, therefore, cannot capture the dynamic behavior of cyber-risk. Addressing this need, the main contribution of this work consists in proposing a modeling framework for optimal cybersecurity investment in a dynamic stochastic setup, allowing for investment policies which respond in real time to randomly occurring cyberattacks. Our work therefore contributes both to cyber-risk modeling and to cyber-risk management, as categorized in the recent survey by [13]. Moreover, our stochastic modeling framework takes into account the empirically relevant feature of temporally clustered cyberattacks, as discussed in the next subsection.
@@ 1.1 Modeling Cyberattacks with Hawkes Processes
Figure 1: One simulated trajectory of $N$ (top) and $\lambda$ (bottom), for $\alpha=27$, $\lambda_{0}=27$, $\xi=15$, $\beta=9$.
@@ 1.3 Structure of the Paper
In Section 2, we recall the original Gordon-Loeb model and introduce our continuous-time stochastic extension. In Section 3, we formulate the cybersecurity investment problem and characterize the optimal policy. Section 4 details the model parameters and the numerical methods used in our analysis. Section 5 presents the results of our numerical analysis and discusses their practical implications for cyber-risk management. Section 6 concludes.
@@ 2.1 The Gordon-Loeb Model
  • $p\ {\in[0,1]}$: the probability that a cyberattack occurs;
  • $v\ {\in[0,1]}$: the probability that a cyberattack successfully breaches the information set (vulnerability);
@@
Without any cybersecurity investment, the expected loss is $vp\ell$. To mitigate its vulnerability, the entity may invest an amount $z\geq 0$ in cybersecurity. The effectiveness of this investment is measured by a security breach probability function $S(z,v)$, which represents the probability that an attack successfully breaches the information set, given investment level $z$ and initial vulnerability $v$. The resulting expected loss is thus $S(z,v)p\ell$. Gordon and Loeb require the function $S$ to satisfy the properties listed in the following assumption.
@@
    $S(0,v)=v$, i.e., in the absence of any investment, the information set retains its baseline vulnerability;
@@ 2.2 A Continuous-Time Model Driven by a Hawkes Process
The cumulative loss incurred over a planning horizon $[0,T]$, in the absence of any cybersecurity investment, is given by:
@@
In our dynamic model, the entity can react to the evolving threat environment by investing in cybersecurity, in order to mitigate its vulnerability to cyberattacks. Investment occurs continuously throughout the planning horizon $[0,T]$ and is described by a non-negative investment rate process $(z_{t})_{t\in[0,T]}$. For each $t<T$, the quantity $z_{t}$ represents the increase in the level of cybersecurity over the infinitesimal time interval $[t,t+\mathrm{d}t]$ and is assumed to be chosen based on the information set $\mathcal{F}_{t}$ available to the decision maker at time $t$.
@@
In the terminology of stochastic processes, the condition that the investment rate $z_{t}$ must be chosen on the basis of $\mathcal{F}_{t}$ corresponds to requiring the process $(z_{t})_{t\in[0,T]}$ to be adapted to the filtration generated by $N$. We point out that, in our setup, the outcomes of previous attacks (i.e., whether breaches have occurred or not) do not carry any relevant informational content for decision making, as they do not affect the dynamics of future attack arrivals.
@@
$\mathrm{d}H_{t}=(z_{t}-\rho H_{t})\mathrm{d}t,\qquad H_{0}\geq 0.$
@@
$S(H_{\tau_{i}},v),$ (7)
where $H_{\tau_{i}}$ is given by (6) evaluated at $t=\tau_{i}$ and $S$ is a security breach probability function satisfying Assumption (A), as in the original Gordon-Loeb model. Hence, the probability that the $i$-th attack successfully breaches the IT system depends on the cybersecurity level $H_{\tau_{i}}$ reached at the attack’s time $\tau_{i}$. In turn, $H_{\tau_{i}}$ is determined by the investment realized over the time period $[0,\tau_{i}]$, taking into account technological obsolescence. If the $i$-th attack breaches the IT system, then the entity incurs into a loss of $\eta_{i}$, otherwise the attack is blocked and the entity does not suffer any loss at time $\tau_{i}$.
@@
$L_{T}^{z}:=\sum_{i=1}^{N_{T}}\eta_{i}B_{i}^{S(H_{\tau_{i}},v)},$ (8)
where $(B_{i}^{S(H_{\tau_{i}},v)})_{i\in\mathbb{N}^{*}}$ is a family of random variables taking values in $\{0,1\}$ and satisfying the following assumption.
@@
$\mathbb{P}\Bigl{(}B_{i}^{S(H_{\tau_{i}},v)}=1\Big{|}\mathcal{F}_{T}\Bigr{)}=S(H_{\tau_{i}},v),\qquad\text{ for all }i\in\mathbb{N}^{*},$
where $(H_{t})_{t\in[0,T]}$ is determined by $(z_{t})_{t\in[0,T]}$ as in (6). Moreover, for each $i\in\mathbb{N}^{*}$, the random variables $B_{i}^{S(H_{\tau_{i}},v)}$ and $\eta_{i}$ are conditionally independent given $\mathcal{F}_{T}$.
@@
The cumulative loss process $(L^{z}_{t})_{t\in[0,T]}$ defined as in (8) constitutes a marked Hawkes process, in the terminology of point processes (see [24]). In our modeling framework, the marks (losses) are endogenous and depend on the dynamically evolving cybersecurity level $(H_{t})_{t\in[0,T]}$.
@@
$\mathbb{E}[L_{T}^{z}]$ $=\bar{\eta}\,\mathbb{E}\left[\int_{0}^{T}S(H_{t},v)\lambda_{t}\mathrm{d}t\right].$
@@
$\mathbb{E}[L_{T}^{0}-L_{T}^{z}]$ $=\bar{\eta}\,\mathbb{E}\left[\int_{0}^{T}\bigl{(}v-S(H_{t},v)\bigr{)}\lambda_{t}\mathrm{d}t\right].$ (9)
@@
$\mathbb{E}[L^{z}_{T}]$ $=\mathbb{E}\left[\sum_{i=1}^{N_{T}}\eta_{i}B_{i}^{S(H_{\tau_{i}},v)}\right]$
$=\mathbb{E}\left[\sum_{i=1}^{N_{T}}\mathbb{E}\left[\eta_{i}B_{i}^{S(H_{\tau_{i}},v)}\Big{|}\mathcal{F}_{T}\right]\right]$
$=\bar{\eta}\,\mathbb{E}\left[\sum_{i=1}^{N_{T}}S(H_{\tau_{i}},v)\right]$
$=\bar{\eta}\,\mathbb{E}\left[\int_{0}^{T}S(H_{t},v)\mathrm{d}N_{t}\right]$
$=\bar{\eta}\,\mathbb{E}\left[\int_{0}^{T}S(H_{t},v)\lambda_{t}\mathrm{d}t\right],$
where the last step follows by definition of intensity (see, e.g., [24, Definition II.D7]), together with the continuity of the process $H$. The first equation in the statement of the proposition follows as a special case by taking $z\equiv 0$. ∎
@@ 3 Optimal Cybersecurity Investment
To ensure the well-posedness of the optimization problem, we constrain the admissible investment policies to a suitably defined admissible set $\mathcal{Z}$ [^5] We point out that, as a direct consequence of the Cauchy-Schwarz inequality, the integral in (6) is always well-defined for every process $z\in\mathcal{Z}$.
@@
The admissible set $\mathcal{Z}$ is defined as the set of all non-negative processes $(z_{t})_{t\in[0,T]}$ such that $\mathbb{E}[\int_{0}^{T}z_{t}^{2}\mathrm{d}t]<\infty$.
@@
$\sup_{z\in\mathcal{Z}}\mathbb{E}\left[L_{T}^{0}-L_{T}^{z}-\int_{0}^{T}\left(\delta z_{t}+\frac{\gamma}{2}z_{t}^{2}\right)\mathrm{d}t+U(H_{T})\right],$ (10)
where $L^{0}_{T}$ and $L^{z}_{T}$ are defined in (5) and (8), respectively, and the state variables $\lambda$ and $H$ satisfy the dynamics
@@
$\mathrm{d}H_{t}$ $=(z_{t}-\rho H_{t})\mathrm{d}t.$ (12)
In the objective functional (10), the term $\mathbb{E}[L_{T}^{0}-L_{T}^{z}]$ represents the reduction in the expected losses due to the investment in cybersecurity. Differently from (3), we consider in (10) a non-linear cost function $z\mapsto\delta z+\gamma z^{2}/2$, for $\delta,\gamma>0$. The non-linearity penalizes irregular or highly concentrated investment strategies, reflecting real-world constraints and incentivizing smoother, more sustained cybersecurity efforts (e.g., continuous IT updates versus abrupt large-scale interventions). The term $U(H_{T})$ accounts for the residual utility of the cybersecurity level reached at the end of the planning horizon. This accounts for the fact that cybersecurity investment carries a long-term benefit, since the entity does not cease to exist after the planning horizon. As usual, the function $U:\mathbb{R}_{+}\to\mathbb{R}$ is assumed to be an increasing and concave utility function.
@@
$\sup_{z\in\mathcal{Z}}\mathbb{E}\left[\int_{0}^{T}\left(\bar{\eta}\bigl{(}v-S(H_{t},v)\bigr{)}\lambda_{t}-z_{t}-\frac{\gamma}{2}z_{t}^{2}\right)\mathrm{d}t+U(H_{T})\right].$ (13)
Problem (13) is a bi-dimensional stochastic optimal control problem, where the stochastic intensity process $(\lambda_{t})_{t\in[0,T]}$ acts as an additional state variable beyond the controlled process $(H_{t})_{t\in[0,T]}$. Due to the Markovian structure of the system, dynamic programming techniques can be applied for the solution of (13). To this end, we first introduce the following notation, for any $(t,\lambda,h)\in[0,T]\times(0,\infty)\times\mathbb{R}_{+}$:
  • $H_{s}^{t,h,z}:=he^{-\rho(s-t)}+\int_{t}^{s}e^{-\rho(s-v)}z_{v}\mathrm{d}v,$ (14)
    for all $s\in[t,T]$, representing the cybersecurity level reached at time $s$ when starting from level $H_{t}=h$ at time $t$ and investing according to a process $z\in\mathcal{Z}$;
  • $\lambda_{s}^{t,\lambda}:=\alpha+(\lambda-\alpha)e^{-\xi(s-t)}+\beta\sum_{i=N_{t}+1}^{N_{s}}e^{-\xi(s-\tau_{i})},$
    for all $s\in[t,T]$, representing the stochastic intensity at time $s$ when starting from value $\lambda_{t}=\lambda$ at time $t$.
For any stopping time $\tau$ taking values in $[0,T]$, we denote by $\mathcal{Z}_{\tau}$ the admissible set $\mathcal{Z}$ restricted to the stochastic time interval $[\tau,T]$.
@@
$J(t,\lambda,h;z)$ $:=\mathbb{E}\left[\int_{t}^{T}\bar{\eta}\bigl{(}v-S(H_{s}^{t,h,z},v)\bigr{)}\lambda_{s}^{t,\lambda}\mathrm{d}s\right.$
$\qquad\quad\left.-\int_{t}^{T}\left(z_{s}+\frac{\gamma}{2}z_{s}^{2}\right)\mathrm{d}s+U(H_{T}^{t,h,z})\right].$
Consequently, the value function associated to the stochastic optimal control problem (13) is given by
@@
  • for every $(t,h)\in[0,T)\times\mathbb{R}_{+}$, the map $\lambda\mapsto V(t,\lambda,h)$ is strictly increasing and convex: the benefit of cybersecurity investment is greater in the presence of a greater risk of cyberattacks, with an increasing marginal benefit of cybersecurity investment;
@@
We now proceed to characterize the value function $V$ as the solution to a Hamilton-Jacobi-Bellman (HJB) partial integro-differential equation (PIDE). To this end, we introduce the following standard assumption.
Assumption (D).
The dynamic programming principle holds: for all $(t,\lambda,h)$ in $[0,T)\times(0,\infty)\times\mathbb{R}_{+}$ and for every stopping time $\tau$ taking values in $[t,T]$, it holds that
$V(t,\lambda,h)$ $=\sup_{z\in\mathcal{Z}_{t}}\mathbb{E}\left[\int_{t}^{\tau}\bar{\eta}\bigl{(}v-S(H_{s}^{t,h,z},v)\bigr{)}\lambda_{s}^{t,\lambda}\mathrm{d}s\right.$
$\qquad\quad\left.-\int_{t}^{\tau}\left(z_{s}+\frac{\gamma}{2}z_{s}^{2}\right)\mathrm{d}s+V(\tau,\lambda_{\tau}^{t,\lambda},H_{\tau}^{t,h,z})\right].$
@@
Suppose that Assumption (D) holds. Assume furthermore that the value function $V$ defined in (15) is of class $\mathcal{C}^{1,1,1}$ (i.e., continuously differentiable in all its arguments). Then, the function $V$ solves the following HJB-PIDE: [^6] For brevity of notation, in the statement and in the proof of this theorem, we omit to denote explicitly the dependence of $V$ on its arguments $(t,\lambda,h)$.
@@
In view of Assumption (D) and the assumption that $V$ is of class $\mathcal{C}^{1,1,1}$, standard arguments based on Itô’s formula together with (11) and (12) imply that $V$ satisfies the following HJB equation (see, e.g., [25, Section 5.2]):
00 0 $=\sup_{z\geq 0}\left(\frac{\partial V}{\partial t}+\xi(\alpha-\lambda)\frac{\partial V}{\partial\lambda}-\rho h\frac{\partial V}{\partial h}+z\frac{\partial V}{\partial h}\right.$
@@
The expectation $\mathbb{E}[\int_{t}^{T}\lambda_{s}^{t,\lambda}\mathrm{d}s]$ can be computed by a straightforward adaptation of Proposition 1 (compare also with [26, Theorem 3.6]), thus completing the proof. ∎

### v2

@@ 1 Introduction
The problem of optimal cybersecurity investment has been first addressed in the seminal work of Gordon and Loeb [24]. In their model, reviewed in Section 2.1 below, the decision maker can reduce the vulnerability to cyberattacks by investing in cybersecurity. The optimal expenditure in cybersecurity is determined by maximizing the expected net benefit of reducing the breach probability. The Gordon-Loeb model laid the foundations for a rigorous quantitative analysis of cybersecurity investments and has been the subject of numerous extensions and generalizations: we mention here only some studies that are closely related to our context, referring to [21] for a comprehensive overview. The key ingredient of the Gordon-Loeb model is represented by the security breach probability function (see Section 2.1), which has been further analyzed in [34] and [44]. The risk-neutral assumption of the original model [24] has been relaxed to accommodate risk-averse preferences in [45]. The practical applicability of the Gordon-Loeb model for guiding cost-effective cybersecurity investment decisions has been examined in [23] within the National Institute of Standards and Technology (NIST) cybersecurity framework.
@@
An effective cybersecurity investment policy must be adaptive and evolve in response to changing threat environments. As noted by [58], a key feature of cyber-risk is its dynamic nature, due to the rapid technological transformation and the evolution of threat actors. Similarly, [3] emphasize the need for adaptable and responsive cybersecurity policies in order to face the challenge of dynamic cyberattacks. The framework of [35] is based on a deterministic model and, therefore, cannot capture the dynamic behavior of cyber-risk. Addressing this need, the main contribution of this work consists in proposing a modeling framework for optimal cybersecurity investment in a dynamic stochastic setup, allowing for investment policies which respond in real time to randomly occurring cyberattacks. Our work contributes both to cyber-risk modeling and to cyber-risk management, as categorized in the recent survey by [29]. Moreover, our stochastic modeling framework takes into account the empirically relevant feature of temporally clustered cyberattacks, as discussed in the next subsection.
@@ 1.1 Modeling Cyberattacks with Hawkes Processes
Figure 1: One simulated trajectory of NN (top) and $\lambda$ (bottom), for $\alpha=27$, $\lambda_{0}=27$, $\xi=15$, $\beta=9$.
@@ 1.3 Structure of the Paper
In Section 2, we recall the original Gordon-Loeb model and introduce our continuous-time stochastic extension. In Section 3, we formulate the cybersecurity investment problem and characterize the optimal policy. Section 4 details the model parameters and the numerical methods used in our analysis. Section 5 presents the results of our numerical analysis and discusses their practical implications for cyber-risk management. Section 6 discusses the relevance of our study from an insurance perspective, also in relation to the calculation of premia for cyber-insurance contracts. Section 7 concludes.
@@ 2.1 The Gordon-Loeb Model
  • p[0,1]p\ $\in$[0,1]}: the probability that a cyberattack occurs;
  • v[0,1]v\ $\in$[0,1]}: the baseline probability that a cyberattack successfully breaches the information set (vulnerability);
@@
The vulnerability $v$ refers to a baseline level of security, reflecting the cybersecurity measures already in place before any new decision is taken. Without any further investment in cybersecurity, the expected loss is $vp\ell$. To mitigate its vulnerability, the entity may invest an additional amount $z\geq 0$ in cybersecurity. The variable $z$ therefore represents incremental cybersecurity investment beyond the existing baseline protection. In particular, $z=0$ means that no further investment is made, but the system still retains the prior protection embodied in $v$. Throughout the paper, absence of cybersecurity investment will refer to the case $z=0$. The effectiveness of the investment $z$ is measured by a security breach probability function $S(z,v)$, which represents the probability that an attack successfully breaches the information set, given investment level $z$ and the vulnerability $v$. The resulting expected loss is thus $S(z,v)p\ell$. Gordon and Loeb require the function $S$ to satisfy the properties listed in the following assumption.
@@
    $S(0,v)=v$, i.e., in the absence of investment, the information set retains its baseline vulnerability;
@@ 2.2 A Continuous-Time Model Driven by a Hawkes Process
Remark 2.
The assumption that the losses $(\eta_{i})_{i\in\mathbb{N}^{*}}$ are independent of the attack arrival process $N$ can be restrictive from an empirical viewpoint. In practice, both the frequency and the severity of cyber losses can increase during periods of clustered attacks. From a modeling standpoint, this assumption can be relaxed, for instance by allowing the loss distribution to depend on the current intensity, making use of the techniques recently developed in [30]. However, introducing such dependency would significantly complicate the analysis of the optimization problem in Section 3 and requires precise information on the conditional distribution of losses given the attack dynamics. In the absence of sufficient publicly available data on cyber losses, we retain the independence assumption as a tractable and transparent modeling choice.
The cumulative loss incurred over a planning horizon $[0,T]$, in the absence of cybersecurity investment, is given by:
@@
In our dynamic model, the entity can react to the evolving threat environment by investing in cybersecurity, in order to mitigate its vulnerability to cyberattacks. Investment occurs continuously throughout the planning horizon $[0,T]$ and is described by a non-negative investment rate process $(z_{t})_{t\in[0,T]}$. For each $t<T$, the quantity $z_{t}$ represents the increase in the level of cybersecurity over the infinitesimal time interval $t,t+\mathrm{d}t]$ and is assumed to be chosen based on the information set $\mathcal{F}_{t-}$ available to the decision maker just before time $t$.
@@
In the terminology of stochastic processes, the requirement that the investment rate $z_{t}$ must be chosen on the basis of $\mathcal{F}_{t-}$ corresponds to requiring the process $(z_{t})_{t\in[0,T]}$ to be predictable with respect to the filtration generated by $N$. In particular, at any attack time $\tau_{i}$, the value $z_{\tau_{i}}$ is determined by the information $\mathcal{F}_{\tau_{i}-}$: the decision maker may adjust investment after observing an attack, but not in a way that anticipates the attack at the same instant. We point out that, in our setup, the outcomes of previous attacks (i.e., whether an attack has successfully breached the system or not) do not carry any relevant informational content for decision making, as they do not affect the dynamics of future attack arrivals.
@@
$\mathrm{d}H^{z}_{t}=(z_{t}-\rho H^{z}_{t})\mathrm{d}t,\qquad H^{z}_{0}=H_{0}\geq 0.$
@@
$S(H^{z}_{\tau_{i}},v),$ (7)
where $H^{z}_{\tau_{i}}$ is given by (6) evaluated at $t=\tau_{i}$ and $S$ is a security breach probability function satisfying Assumption (A), as in the original Gordon-Loeb model. Hence, the probability that the $i$-th attack successfully breaches the IT system depends on the cybersecurity level $H^{z}_{\tau_{i}}$ reached at the attack’s time $\tau_{i}$. In turn, $H^{z}_{\tau_{i}}$ is determined by the investment realized over the time period $0,\tau_{i}]$, taking into account technological obsolescence. If the $i$-th attack breaches the IT system, then the entity incurs into a loss of $\eta_{i}$, otherwise the attack is blocked and the entity does not suffer any loss at time $\tau_{i}$.
@@
$L_{T}^{z}:=\sum_{i=1}^{N_{T}}\eta_{i}B_{i}^{S(H^{z}_{\tau_{i}},v)},$ (8)
where $(B_{i}^{S(H^{z}_{\tau_{i}},v)})_{i\in\mathbb{N}^{*}}$ is a family of random variables taking values in $\{0,1\}$ and satisfying the following assumption.
@@
$\mathbb{P}\Bigl(B_{i}^{S(H^{z}_{\tau_{i}},v)}=1\Big|\mathcal{F}_{T}\Bigr)=S(H^{z}_{\tau_{i}},v),\qquad\text{ for all }i\in\mathbb{N}^{*},$
where $(H^{z}_{t})_{t\in[0,T]}$ is determined by $(z_{t})_{t\in[0,T]}$ as in (6). Moreover, for each $i\in\mathbb{N}^{*}$, the random variables $B_{i}^{S(H^{z}_{\tau_{i}},v)}$ and $\eta_{i}$ are conditionally independent given $\mathcal{F}_{T}$.
@@
The cumulative loss process $(L^{z}_{t})_{t\in[0,T]}$ defined as in (8) constitutes a marked Hawkes process, in the terminology of point processes (see [13]). In our modeling framework, the marks (losses) are endogenous and depend on the dynamically evolving cybersecurity level $(H^{z}_{t})_{t\in[0,T]}$.
@@
$\mathbb{E}[L_{T}^{z}]$ $=\bar{\eta}\,\mathbb{E}\left[\int_{0}^{T}S(H^{z}_{t},v)\lambda_{t}\mathrm{d}t\right].$
@@
$\mathbb{E}[L_{T}^{0}-L_{T}^{z}]$ $=\bar{\eta}\,\mathbb{E}\left[\int_{0}^{T}\bigl(v-S(H^{z}_{t},v)\bigr)\lambda_{t}\mathrm{d}t\right].$ (9)
@@
$\mathbb{E}[L^{z}_{T}]$ $=\mathbb{E}\left[\sum_{i=1}^{N_{T}}\eta_{i}B_{i}^{S(H^{z}_{\tau_{i}},v)}\right]$
$=\mathbb{E}\left[\sum_{i=1}^{N_{T}}\mathbb{E}\left[\eta_{i}B_{i}^{S(H^{z}_{\tau_{i}},v)}\Big|\mathcal{F}_{T}\right]\right]$
$=\bar{\eta}\,\mathbb{E}\left[\sum_{i=1}^{N_{T}}S(H^{z}_{\tau_{i}},v)\right]$
$=\bar{\eta}\,\mathbb{E}\left[\int_{0}^{T}S(H^{z}_{t},v)\mathrm{d}N_{t}\right]$
$=\bar{\eta}\,\mathbb{E}\left[\int_{0}^{T}S(H^{z}_{t},v)\lambda_{t}\mathrm{d}t\right],$
where the last step follows by definition of intensity (see, e.g., [13, Definition II.D7]), together with the continuity of the process $H^{z}$, which implies that $H^{z}_{\tau_{i}}$ coincides with the left-limit $H^{z}_{\tau_{i}-}$, for all $i\in\mathbb{N}^{*}$. The first equation in the statement of the proposition follows as a special case by taking $z\equiv 0$. ∎
@@ 3 Optimal Cybersecurity Investment
To ensure the well-posedness of the optimization problem, we constrain the admissible investment policies to a suitably defined admissible set $\mathcal{Z}$. [^5] We point out that, as a direct consequence of the Cauchy-Schwarz inequality, the integral in (6) is always well-defined for every process $z\in\mathcal{Z}$. In line with Remark 3, in the following definition we restrict our attention to processes that are predictable with respect to the filtration generated by $N$.
@@
The admissible set $\mathcal{Z}$ is defined as the set of all non-negative predictable processes $(z_{t})_{t\in[0,T]}$ such that $\mathbb{E}[\int_{0}^{T}z_{t}^{2}\mathrm{d}t]<\infty$.
@@
$\sup_{z\in\mathcal{Z}}\mathbb{E}\left[L_{T}^{0}-L_{T}^{z}-\int_{0}^{T}\left(\delta z_{t}+\frac{\gamma}{2}z_{t}^{2}\right)\mathrm{d}t+U(H^{z}_{T})\right],$ (10)
where $L^{0}_{T}$ and $L^{z}_{T}$ are defined in (5) and (8), respectively, and the state variables $\lambda$ and $H^{z}$ satisfy the dynamics
@@
$\mathrm{d}H^{z}_{t}$ $=(z_{t}-\rho H^{z}_{t})\mathrm{d}t.$ (12)
In the objective functional (10), the term $\mathbb{E}[L_{T}^{0}-L_{T}^{z}]$ represents the reduction in the expected losses due to the investment in cybersecurity. Differently from (3), we consider in (10) a non-linear cost function $z\mapsto\delta z+\gamma z^{2}/2$, for $\delta,\gamma>0$. The non-linearity penalizes irregular or highly concentrated investment strategies, reflecting real-world constraints and incentivizing smoother, more sustained cybersecurity efforts (e.g., continuous IT updates versus abrupt large-scale interventions). The term $U(H^{z}_{T})$ accounts for the residual utility of the cybersecurity level reached at the end of the planning horizon. This accounts for the fact that cybersecurity investment carries a long-term benefit, since the entity does not cease to exist after the planning horizon. As usual, the function $U:\mathbb{R}_{+}\to\mathbb{R}$ is assumed to be an increasing and concave utility function.
@@
$\sup_{z\in\mathcal{Z}}\mathbb{E}\left[\int_{0}^{T}\left(\bar{\eta}\bigl(v-S(H^{z}_{t},v)\bigr)\lambda_{t}-z_{t}-\frac{\gamma}{2}z_{t}^{2}\right)\mathrm{d}t+U(H^{z}_{T})\right].$ (13)
Remark 6.
The objective functional (10) is linear in the losses $(\eta_{i})_{i\in\mathbb{N}^{*}}$. More specifically, formula (13) shows that the loss distribution enters into the optimization problem only through its first moment. In turn, this implies that the optimal policy derived in Theorem 1 below depends only on $\bar{\eta}$ and not on higher-order moments of the loss distribution. Under nonlinear objective functionals, higher moments and the tail behavior of the loss distribution would affect the optimal policy.
Problem (13) is a bi-dimensional stochastic optimal control problem, where the stochastic intensity process $(\lambda_{t})_{t\in[0,T]}$ acts as an additional state variable beyond the controlled process $(H^{z}_{t})_{t\in[0,T]}$. Due to the Markovian structure of the system, dynamic programming techniques can be applied for the solution of (13). To this end, we first introduce the following notation, for any $(t,\lambda,h)\in[0,T]\times(0,\infty)\times\mathbb{R}_{+}$:
  • for all $s\in[t,T]$,
    $H_{s}^{t,h;z}:=he^{-\rho(s-t)}+\int_{t}^{s}e^{-\rho(s-v)}z_{v}\mathrm{d}v$ (14)
    represents the cybersecurity level reached at time $s$ when starting from level $H_{t}=h$ at time $t$ and investing according to $z\in\mathcal{Z}$;
  • for all $s\in[t,T]$,
    $\lambda_{s}^{t,\lambda}:=\alpha+(\lambda-\alpha)e^{-\xi(s-t)}+\beta\sum_{i=N_{t}+1}^{N_{s}}e^{-\xi(s-\tau_{i})}$
    represents the stochastic intensity at time $s$ when starting from value $\lambda_{t}=\lambda$ at time $t$.
For any stopping time $\tau$ taking values in $[0,T]$, we denote by $\mathcal{Z}_{\tau}$ the admissible set $\mathcal{Z}$ restricted to the stochastic time interval $\tau,T]$.
@@
$J(t,\lambda,h;z)$ $:=\mathbb{E}\left[\int_{t}^{T}\bar{\eta}\bigl(v-S(H_{s}^{t,h;z},v)\bigr)\lambda_{s}^{t,\lambda}\mathrm{d}s\right$.
$\qquad\quad\left.-\int_{t}^{T}\left(z_{s}+\frac{\gamma}{2}z_{s}^{2}\right)\mathrm{d}s+U(H_{T}^{t,h;z})\right].$
As a consequence of Definition 1 together with the concavity of the function $U$, the functional $J$ is always well-defined and finite. Consequently, the value function associated to the stochastic optimal control problem (13) is given by
@@
  • for every $(t,h)\in[0,T)\times\mathbb{R}_{+}$, the map $\lambda\mapsto V(t,\lambda,h)$ is strictly increasing: the benefit of cybersecurity investment is greater in the presence of a greater risk of cyberattacks;
@@
The optimization problem formulated in (10) falls within the framework of the optimal control of a piecewise deterministic Markov process, in the sense of [17]. Under the admissibility conditions imposed in Definition 1, the performance functional $J$ is well-posed for every control $z\in\mathcal{Z}_{t}$. As a consequence, the same techniques used in the proof of [51, Theorem 3.3.1] yield that the value function satisfies the dynamic programming principle. More precisely, for all $(t,\lambda,h)$ in $0,T)\times(0,\infty)\times\mathbb{R}_{+}$ and for every stopping time $\tau$ taking values in $[t,T]$, it holds that
$V(t,\lambda,h)$ $=\sup_{z\in\mathcal{Z}_{t}}\mathbb{E}\left[\int_{t}^{\tau}\bar{\eta}\bigl(v-S(H_{s}^{t,h;z},v)\bigr)\lambda_{s}^{t,\lambda}\mathrm{d}s\right$.
$\qquad\quad\left.-\int_{t}^{\tau}\left(z_{s}+\frac{\gamma}{2}z_{s}^{2}\right)\mathrm{d}s+V(\tau,\lambda_{\tau}^{t,\lambda},H_{\tau}^{t,h;z})\right].$
We now proceed to characterize the value function $V$ as the solution to a Hamilton-Jacobi-Bellman (HJB) partial integro-differential equation (PIDE). Further theoretical results on the value function, together with a verification theorem under suitable regularity assumptions, are proved in [50, Section 2.3].
@@
Suppose that the value function $V$ defined in (15) is of class $\mathcal{C}^{1,1,1}$ (i.e., continuously differentiable in all its arguments). Then, the function $V$ solves the following HJB-PIDE: [^6] For brevity of notation, in the statement and in the proof of this theorem, we omit to denote explicitly the dependence of $V$ on its arguments $(t,\lambda,h)$.
@@
In view of the assumption that $V$ is of class $\mathcal{C}^{1,1,1}$, standard arguments based on Itô’s formula together with (11) and (12) imply that $V$ satisfies the following HJB equation (see, e.g., [4, Section 5.2]):
$0$ $=\sup_{z\geq 0}\left(\frac{\partial V}{\partial t}+\xi(\alpha-\lambda)\frac{\partial V}{\partial\lambda}-\rho h\frac{\partial V}{\partial h}+z\frac{\partial V}{\partial h}\right$.
@@
The expectation $\mathbb{E}[\int_{t}^{T}\lambda_{s}^{t,\lambda}\mathrm{d}s]$ can be computed by a straightforward adaptation of Proposition 1 (compare also with [15, Theorem 3.6]), thus completing the proof. ∎